OpenStudy (anonymous):

Which of the following are exponential functions? (Check all that apply.) A. y = 1 + x B. y = 1.2x - 2 C. y = 4.2(5)x D. y = 5x - 72 Select all that apply

OpenStudy (anonymous):

none...

sam (.sam.):

No

sam (.sam.):

I mean none, lol

OpenStudy (anonymous):

I chose A,D..

Directrix (directrix):

Question: Does "C. y = 4.2(5)x" mean y = 4.2 * 5 ^x ? If so, C is the equation of exponential function.

Directrix (directrix):

Option A is a line. Option D is a line as it. If D is meant to be y = 5^x - 72, then D is exponential.